# Fairness in Serving Large Language Models

### v1

@@ 4 Achieving Fairness
In this section, we present our algorithm VTC with proved fairness properties in Section˜4.1, and show its generalization for customized service measurement in Section˜4.2.
@@ 4.1 Virtual Token Counter (VTC)
Based on insights from prior discussions, we’ve identified three key challenges inherent in Large Language Model (LLM) serving that hinder the effective adaptation of existing algorithms for delivering approximately fair LLM service: (1) The cost per token can vary, i.e., the service depends on both the number of input and output tokens, (2) The server has variable token-rate capacity, and (3) the output length is unpredictable.
To address these three challenges, we propose the Virtual Token Counter (VTC), a mechanism for achieving fair sharing in LLM Serving (Algorithm˜2). To quantify the service received by a client we use the weighted number of tokens metric, as described in Section˜3.1. We discuss the generalization to other metrics in Section˜4.2.
@@
Algorithm˜2 shows how VTC can be implemented in the continuous batching framework described in Section˜2.1. It maintains a virtual counter for each client, denoted as $\{c_{i}\}$. The counters are initialized as 0 (line 2). The program runs with two parallel streams.
The execution stream is the control loop of an execution engine that implements continuous batching. Line 17 controls the frequency of adding a minibatch $B_{new}$ of new requests into the running batch $B$. Commonly, the server will add a new minibatch after several decoding steps. The minibatch $B_{new}$ is constructed by iteratively selecting the request from the client with the smallest virtual counter (lines 20-25). The counters will be updated when adding new requests according to the service invoked by the input tokens (line 23). After each decode step (line 28), $\{c_{i}\}$ will be updated immediately according to the service invoked by the newly generated output tokens (line 29).
@@
1:request trace, input token weight $w_{p}$, output token weight $w_{q}$, upper bound from Equation˜2 denoted as $U$.
@@
23:     if $r$ cannot fit in the memory then Break
24:     $c_{k}\leftarrow c_{k}+w_{p}\cdot input\_length(r)$
25:     $B_{new}\leftarrow B_{new}+r$
26:     $Q\leftarrow Q-r$
[…]
28:   $B\leftarrow B+B_{new}$
[…]
30:  $c_{i}\leftarrow c_{i}+w_{q}\cdot|\{r\mid client(r)=i,r\in B\}|$
[…]
@@ 4.1.1 Fairness for backlogged clients in VTC
We adapt the traditional definition of fairness for backlogged clients in the network to our scenario as follows.
@@
In the rest of the paper, as in Algorithm˜2, we let $Q$ denote the set of requests in the waiting queue. We abuse the notation of $i\in Q$ for a client $i$ to indicate there exists $r\in Q$, such that $r$ is a request from client $i$. Let $L_{input}$ and $L_{output}$ be the maximum number of input and output tokens in a request. Let $M$ be the maximum number of tokens of a client that can be fitted in a running batch. Next, we start with some lemmas to prove a bound for Definition 4.2 in Theorem˜4.4. The missing proofs are in Appendix˜A.
@@
The following invariant holds at any time in Algorithm˜2 when $Q\neq\emptyset$:
@@
For any clients $f$ and $g$, for any time interval $[t_{1},t_{2})$ in which $f$ and $g$ are backlogged, Algorithm˜2 guarantees  [^4] The service of a served request incurred by pre-filling (service for input tokens) is counted at the time when the request is added to the running batch (line 23 in Algorithm 2), rather than the time when prefill is finished. This is because we want to count the input tokens immediately to avoid selecting all the same $k$ at line 20 in Algorithm 2 for $B_{new}$.
@@
For any $f$, if $f$ is backlogged during time $t_{1}$ to $t_{2}$, we have $W_{f}(t_{1},t_{2})=c_{f}^{(t_{2})}-c_{f}^{(t_{1})}$. This is because the line 7 will not be reached for client $f$ during $t_{1}$ to $t_{2}$, and the $c_{f}$ keeps increasing during $t_{1}$ to $t_{2}$ by adding $w_{p}$ product the number of served input tokens and $w_{q}$ product the number of served output tokens. By Lemma 4.3, from Equation˜2, we have
$|\displaystyle|W_{f}(t_{1},t_{2})-W_{g}(t_{1},t_{2})|$$\leq|c_{f}^{(t_{1})}-c_{g}^{(t_{1})}|+|c_{f}^{(t_{2})}-c_{g}^{(t_{2})}|$
@@
Line 13 can be modified to take any value between $\min\{c_{i}|\exists r^{\prime}\in Q,client(r^{\prime})=i\}$ and $\max\{c_{i}|\exists r^{\prime}\in Q,client(r^{\prime})=i\}$. The proof of Theorem˜4.4 should still hold.
@@
To tighten the bound in Theorem˜4.4, we can restrict the memory usage for each client in the running batch. This might compromise the work-conserving property, as we will demonstrate in Theorem˜4.8. Therefore, there is a trade-off between achieving a better fairness bound and maintaining work conservation.
We also prove in the next theorem that the bound in Theorem˜4.4 is tight within a factor of 222 for a family of work-conserving schedulers. We say a scheduler is work-conserving if it stops adding requests to a partially-filled minibatch (line 21 in Algorithm˜2) only when it runs out of memory [^5] Different implementation may have different criteria of “not enough memory”. This can only be achieved heuristically because the number of output tokens is unknown before it finishes. but not for fairness reasons.
@@
As we mentioned before, output tokens are more expensive than input tokens, so normally we have $w_{q}>w_{p}$. Therefore the right-hand side of the inequality in Theorem˜4.4 is $2w_{q}\cdot M$, which is 2×2$\times$2 × of the lower bound in Theorem˜4.8.
@@ 4.1.2 Fairness for non-backlogged clients in VTC
In this subsection, we discuss item 2 in Section˜3.2. A backlogged client will not receive less service than another client. This can be reflected in the following theorem. The missing proofs are in Appendix˜A.
@@
Here $U$ is the upper bound from Equation˜2.
In addition to that, clients who send requests constantly less than their share should have their requests serviced nearly instantly. This property intuitively can be implied by the first item in Section˜3.2, as if a low-rate client cannot be served on time, it becomes backlogged, which requires the same level of service with backlogged clients. We formally prove this property to offer a fairness assurance for clients who are not overloaded in Theorem˜4.13. This intuitively acts as a safeguard against misbehaving clients [7].
We start with Theorem˜4.11 discussing the aspect of latency bounds. Intuitively, if a client is not backlogged and has no requests running, the next request from it will be processed within a latency bound that is independent of the request rate of other clients.
@@
The bound in Theorem˜4.11 is irrelevant to the request rate of others, giving an upper bound for latency against ill-behavior clients.
[…]
@@
Assume a client $f$ is not backlogged at time $t_{1}$ and for any time interval $[t,t_{2}),t_{1}\leq t<t_{2}$, $f$ has requested services less than $\frac{T(t,t_{2})}{n(t,t_{2})}-5U$, where $T(t,t_{2})$ is the total services received for all clients during the interval $[t,t_{2})$, $n(t,t_{2})$ is the number of clients that have requested services during the interval, and $U$ is the upper bound from Equation˜2.
@@ 4.2 Adapt to Different Fairness Criteria
Algorithm˜2 is designed for fairness with the service function $W(t_{1},t_{2})$ as a linear combination of the number of processed input tokens and the number of generated tokens. For a different definition of $W(t_{1},t_{2})$, Algorithm˜2 can be easily modified to update the counter according to the other definitions described in Section˜3.1.
For example, let $n_{in}$ and $n_{out}$ be the number of processed input tokens and generated tokens, respectively. Assume we are pursuing fairness with $h(n_{in},n_{out})$ as the service measurement for some function $h$. Line 23 will be changed to
$c_{k}\leftarrow c_{k}+h(l_{in}^{r},0).$
Line 29 will be changed to
$c_{i}\leftarrow c_{i}+\sum_{r\mid client(r)=i,r\in B}\left(h(l_{in}^{r},n_{out}^{r})-h(l_{in}^{r},n_{out}^{r}-1)\right).$
Here $l_{in}^{r}$ denotes the input length of request $r$, and $n_{out}^{r}$ denotes the number of generated tokens of request $r$.
The fairness bound will also be changed according to $h(n_{in},n_{out})$. Under the assumption that output tokens are more expensive than input tokens, the bound will become the maximum value of aggregated $h(\cdot,\cdot)$ for a set of requests that can be fitted in one running batch.
@@ 5 Evaluations
In this section, we evaluate VTC against other alternatives under different workloads. The results confirm the fairness properties introduced in Section˜3 of VTC, and show that all other alternatives will fail in at least one workload.
@@ 5.1 Setup
Implementation We implement our VTC and other baseline schedulers in S-LoRA[39], a system that serves a large amount of LoRA adapters concurrently. Its backbone is a general serving system adapted from LightLLM [25]. It includes the implementation of continuous batching [46] and PagedAttention [20]666with block size equals 1.. Our VTC scheduler is built on top of those two techniques. Fairness can be considered among general clients, and our experiments are done in this way. But we would like to note that fairness also could be taken into consideration among adapters, especially under the scenario of personalization that uses one adapter per customer, which originally motivated this paper.
@@
  • Request per minute (RPM): This method limits the maximum number of requests that a client can make to the server within a one-minute timeframe. The definition of service corresponds to Section˜3. When a client exceeds this limit, subsequent requests are blocked until the limit resets at the start of the next minute.
@@
We run Llama-2-7b on A10G (24GB), using the memory pool for the KV cache with size 10000777There are in total 10000 tokens for KV cache that can be stored on GPU.. We use various workloads to demonstrate different aspects of fairness, and compare VTC with other baselines. The detailed results are in Section˜5.2. We start with synthetic workloads to give a clear message for fairness properties.

### v2

@@ 4 Achieving Fairness
In this section, we present our algorithm VTC with proved fairness properties in Section 4.1, and show its generalization for customized service measurement in Section 4.2. Variants of VTC, including weighted VTC and VTC with length prediction, are introduced in Section 4.3 and Section 4.4.
@@ 4.1 Virtual Token Counter (VTC)
Based on insights from prior discussions, we’ve identified key challenges inherent in large language model (LLM) serving that hinder direct adaptation of existing algorithms to deliver approximately fair LLM service. We then propose the Virtual Token Counter (VTC), a mechanism for achieving fair sharing in LLM Serving (Algorithm 2). To quantify the service received by a client we use the weighted number of tokens metric, as described in Section 3.1. We discuss the generalization to other metrics in Section 4.2.
@@
1:request trace, input token weight $w_{p}$, output token weight $w_{q}$, upper bound from Equation 2 denoted as $U$.
@@
23:              if $r$ cannot fit in the memory then
24:                  Break
25:              $c_{k}\leftarrow c_{k}+w_{p}\cdot input\_length(r)$
26:              $B_{new}\leftarrow B_{new}+r$
27:              $Q\leftarrow Q-r$
28:         forward_prefill($B_{new}$)
29:         $B\leftarrow B+B_{new}$
30:     forward_decode($B$)
31:     $c_{i}\leftarrow c_{i}+w_{q}\cdot|\{r\mid client(r)=i,r\in B\}|$
32:     $B\leftarrow$ filter_finished_requests($B$)
Algorithm 2 shows how VTC can be implemented in the continuous batching framework described in Section 2.1. A more general integration for VTC is described in Section C.1. It maintains a virtual counter for each client, denoted as $\{c_{i}\}$. The counters are initialized as 00 (line 2). The program runs with two parallel streams.
The execution stream is the control loop of an execution engine that implements continuous batching. Line 17 controls the frequency of adding a minibatch $B_{new}$ of new requests into the running batch $B$. Commonly, the server will add a new minibatch after several decoding steps. The minibatch $B_{new}$ is constructed by iteratively selecting the request from the client with the smallest virtual counter (lines 20-26). The counters will be updated when adding new requests according to the service invoked by the input tokens (line 24). After each decoding step (line 29), $\{c_{i}\}$ will be updated immediately according to the service invoked by the newly generated output tokens (line 30).
@@ 4.1.1 Fairness for backlogged clients in VTC
We adapt the traditional definition of fairness for backlogged clients in the network to our scenario. The following definition formally defined the item 1 introduced in Section 3.2, that for any interval, and any two continuously backlogged clients during the time interval, the difference of their received service should be bounded by a value that is independent of the interval length.
@@
In the rest of the paper, as in Algorithm 2, we let $Q$ denote the set of requests in the waiting queue. We abuse the notation of $i\in Q$ for a client $i$ to indicate there exists $r\in Q$, such that $r$ is a request from client $i$. Let $L_{input}$ and $L_{output}$ be the maximum number of input and output tokens in a request. Let $M$ be the maximum number of tokens that can be fitted in a running batch. Lemma 4.3 reflects the core design of Algorithm 2, that the virtual counters for active clients are chasing each other to ensure their maximum difference is bounded. The missing proof for Lemma 4.3 and all following theorems are in Appendix A.
@@
The following invariant holds at any time in Algorithm 2 when $Q\neq\emptyset$:
@@
We then introduce our main theorem which provides a bound for Definition 4.2.
@@
For any clients $f$ and $g$, for any time interval $[t_{1},t_{2})$ in which $f$ and $g$ are backlogged, Algorithm 2 guarantees  [^5] The service of a served request incurred by pre-filling (service for input tokens) is counted at the time when the request is added to the running batch (line 24 in Algorithm 2), rather than the time when prefill is finished. This is because we want to count the input tokens immediately to avoid selecting all the same $k$ at line 20 in Algorithm 2 for $B_{new}$.
@@
For any $f$, if $f$ is backlogged during time $t_{1}$ to $t_{2}$, we have $W_{f}(t_{1},t_{2})=c_{f}^{(t_{2})}-c_{f}^{(t_{1})}$. This is because the line 7 will not be reached for client $f$ during $t_{1}$ to $t_{2}$, and the $c_{f}$ keeps increasing during $t_{1}$ to $t_{2}$ by adding $w_{p}$ product the number of served input tokens and $w_{q}$ product the number of served output tokens. By Lemma 4.3, from Equation 2, we have
$|W_{f}(t_{1},t_{2})-W_{g}(t_{1},t_{2})|$ $\leq|c_{f}^{(t_{1})}-c_{g}^{(t_{1})}|+|c_{f}^{(t_{2})}-c_{g}^{(t_{2})}|$
@@
Line 13 can be modified to take any value between $\min\{c_{i}|\exists r^{\prime}\in Q,client(r^{\prime})=i\}$ and $\max\{c_{i}|\exists r^{\prime}\in Q,client(r^{\prime})=i\}$. The proof of Theorem 4.4 should still hold.
@@
To tighten the bound in Theorem 4.4, we can restrict the memory usage for each client in the running batch. This might compromise the work-conserving property, as we will demonstrate in Theorem 4.8. Therefore, there is a trade-off between achieving a better fairness bound and maintaining work conservation. Heuristically, predicting the request length in advance could result in a smaller discrepancy, as detailed in Section 4.4. Additionally, preemption is another method to achieve smaller differences, discussed in Section C.3.
We also prove in the next theorem that the bound in Theorem 4.4 is tight within a factor of $2$ for a family of work-conserving schedulers. We say a scheduler is work-conserving if it stops adding requests to a partially-filled minibatch (line 22 in Algorithm 2) only when it runs out of memory [^6] Different implementation may have different criteria of “not enough memory”. This can only be achieved heuristically because the number of output tokens is unknown before it finishes. but not for fairness reasons.
@@
As we mentioned before, output tokens are more expensive than input tokens, so normally we have $w_{q}>w_{p}$. Therefore the right-hand side of the inequality in Theorem 4.4 is $2w_{q}\cdot M$, which is 2×2$\times$2 × of the lower bound in Theorem 4.8.
@@ 4.1.2 Fairness for non-backlogged clients in VTC
In this subsection, we discuss item 2 in Section 3.2. A backlogged client will not receive less service than another client. This can be reflected in the following theorem.
@@
Here $U$ is the upper bound from Equation 2.
In addition to that, clients who send requests constantly less than their share should have their requests serviced nearly instantly. This property intuitively can be implied by the first item in Section 3.2, as if a low-rate client cannot be served on time, it becomes backlogged, which requires the same level of service with backlogged clients. We formally prove this property to offer a fairness assurance for clients who are not overloaded. This intuitively acts as a safeguard against misbehaving clients [10].
We start with Definition 4.10 and Theorem 4.11 discussing the aspect of latency bounds. Intuitively, if a client is not backlogged and has no requests running, the next request from it will be processed within a latency bound that is independent of the request rate of other clients.
@@
The bound in Theorem 4.11 is irrelevant to the request rate of others, giving an upper bound for latency against ill-behavior clients.
The above is about one request not getting delayed. The following theorem shows that during time period $[t_{1},t_{2})$, if there are $n$ active clients sending requests, and client $f$ is sending requests with a rate constantly less than $1/n$ of the server’s capacity (with some constant gap), client $f$ should have all its requests been served.
@@
Assume a client $f$ is not backlogged at time $t_{1}$ and for any time interval $[t,t_{2}),t_{1}\leq t<t_{2}$, $f$ has requested services less than $\frac{T(t,t_{2})}{n(t,t_{2})}-5U$, where $T(t,t_{2})$ is the total services received for all clients during the interval $[t,t_{2})$, $n(t,t_{2})$ is the number of clients that have requested services during the interval, and $U$ is the upper bound from Equation 2.
@@ 4.2 Adapt to Different Fairness Criteria
Algorithm 2 is designed for fairness with the service function $W(t_{1},t_{2})$ as a linear combination of the number of processed input tokens and the number of generated tokens. For a different definition of $W(t_{1},t_{2})$, Algorithm 2 can be easily modified to update the counter according to the other definitions described in Section 3.1.
Assume we aim for fairness using $\sum_{r}h(n_{p}^{r},n_{q}^{r})$ as the metric of service, where $h$ is a specific function. In this context, $r$ indexes the served requests, and $n_{p}^{r},n_{q}^{r}$ represent the number of input and output tokens served for request $r$, respectively. Line 24 will be changed to
$c_{k}\leftarrow c_{k}+h(n_{p}^{r},0).$
Line 30 will be changed to
$c_{i}\leftarrow c_{i}+\sum_{r\mid client(r)=i,r\in B}\left(h(n_{p}^{r},n_{q}^{r})-h(n_{p}^{r},n_{q}^{r}-1)\right).$
The fairness bound will also be changed according to $h(\cdot,\cdot)$. Under the assumption that output tokens are more expensive than input tokens, the bound will become the maximum value of aggregated $h(\cdot,\cdot)$ for a set of requests that can be fitted in one running batch. Algorithm 4 in Section C.1 is the pseudocode of a general VTC framework.
4.3 Weighted VTC
VTC can be applied when clients have tiers. Similar to weighted fair queuing, clients can have different weights to represent their priority in service. If a client $f$ has a weight $w_{1}$, that is twice the weight $w_{2}$ of client $g$, client $f$ is expected to receive twice the service than client $g$. When they are continuously backlogged during the interval $[t_{1},t_{2})$, we want $\left|\frac{W_{f}(t_{1},t_{2})}{w_{1}}-\frac{W_{g}(t_{1},t_{2})}{w_{2}}\right|$ to be bounded instead of $\left|W_{f}(t_{1},t_{2})-W_{g}(t_{1},t_{2})\right|$.
Weighted VTC can be easily implemented by modifying the lines that update the virtual tokens. For example, the line 22 in Algorithm 4 will be changed to
$c_{i}\leftarrow c_{i}+\dfrac{\sum_{r\mid client(r)=i}\left(h(n_{p}^{r},n_{q}^{r})-h(n_{p}^{r(old)},n_{q}^{r(old)})\right)}{w_{i}}.$
Here $c_{i}$ is the virtual counter of client $i$, and $w_{i}$ is its corresponding weight.
4.4 VTC with Length Prediction
As mentioned in Remark 4.7, using VTC with length prediction can heuristically reduce the service discrepancy. In standard VTC, the counters only reflect served tokens. Tokens generated in the future can only be passively added to the counter. This results in a large service discrepancy because requests are overly added due to underestimation of their costs at the time of prompting, leading to the forced serving of over-compensated output tokens. Incorporating a prediction mechanism can help reduce this variance.
The theoretical worst-case scenario won’t change, according to the lower bound proved in Theorem 4.8. But practically, the average-case service discrepancy could be smaller.
The modified pseudocode of VTC with length prediction is described in Algorithm 3 in Section B.3. Intuitively, when a request $r$ is selected, the cost associated with the predicted number of output tokens is immediately added to the virtual counter of the client sending the request. During the actual decoding process, adjustments are made to the virtual counter based on the actual number of output tokens produced. If the actual number of tokens exceeds the prediction, the virtual counter is increased accordingly. Conversely, if fewer tokens are generated than predicted when finished, the virtual counter is reduced. The effectiveness of the length predictor is contingent upon both the workload and the accuracy of predictions, as demonstrated in our evaluations.
@@ 5 Evaluations
In this section, we evaluate VTC against other alternatives under different workloads. The results confirm the fairness properties introduced in Section 3 of VTC, and show that all other alternatives will fail in at least one workload.
@@ 5.1 Setup
Implementation We implement our VTC and other baseline schedulers in S-LoRA[43], a system that serves a large amount of LoRA adapters concurrently. Its backbone is a general serving system adapted from LightLLM [29]. It includes the implementation of continuous batching [50] and PagedAttention [24]777with block size equals 1.. Our VTC scheduler is built on top of those two techniques. Our implementation is elegant and can be implemented as a thin layer on top of the existing scheduler, it contains only about 100 lines of code on top of S-LoRA. The simplicity demonstrates its wide applicability. Fairness can be considered among general clients, and our experiments are done in this way. But we would like to note that fairness also could be taken into consideration among adapters, especially under the scenario of personalization that uses one adapter per customer, which originally motivated this paper.
@@
  • Request per minute (RPM): This method limits the maximum number of requests that a client can make to the server within a one-minute timeframe. The definition of service corresponds to Section 3. When a client exceeds this limit, subsequent requests are blocked until the limit resets at the start of the next minute.
@@
We also benchmark the VTC with length predictions as described below:
  • VTC (predict): This variant of VTC, detailed in Algorithm 3, utilizes the average output length of the last five requests from each client to predict the output length.
  • VTC (oracle): This variant employs a hypothetical output length predictor that achieves 100% accuracy.
@@
We run Llama-2-7b on A10G (24GB), using the memory pool for the KV cache with size 10000888There are in total 10000 tokens for KV cache that can be stored on GPU.. We use various workloads to demonstrate different aspects of fairness, and compare VTC with other baselines. The detailed results are in Section 5.2. We start with synthetic workloads to give a clear message for fairness properties.
@@
Metrics
We apply the weighted number of tokens described in Section 3.1 as the measurement of services in our evaluation. Following OpenAI pricing, we set $w_{p}=1$ and $w_{q}=2$.
  • The service received by client $i$ at time $t$ is measured as $W_{i}(t-T,t+T)$ for a certain $T$.
  • The absolute difference in service between clients is quantified based on accumulated services, represented as $\max_{i,j}|W_{i}(0,t)-W_{j}(0,t)|$.
  • The response time of client $i$ at time $t$ is measured as the average first token latency of the requests sent by client $i$ during the time window $[t-T,t+T]$.
In all settings, we set $T=30$ seconds.
We employ service difference as a quantitative metric to assess the deviation from ideal fairness. A smaller difference in service indicates more equitable scheduling. Formally, the service difference between two clients is defined as the minimum of two values: the difference between their received services, and the difference between the lower service and its corresponding request rate. For example, consider two clients that received services $s_{1}$ and $s_{2}$, such that $s_{1}\leq s_{2}$, and let $r_{1}$ denote the request rate sent from the first client. Then the service difference is defined as $\min(s_{2}-s_{1},|r_{1}-s_{1}|)$.
VTC Variants
The experiments for weighted VTC are presented in Section B.1, demonstrating its capability to serve clients with varying priorities. To illustrate the versatility of the service function beyond the linear model used in our primary analysis, we evaluate a profiled service cost function in Section B.2, which is a quadratic function. Additional experiments on VTC with length prediction are detailed in Section B.3.